OpenStudy (migitmack):

Frank kicks a soccer ball off the ground and in the air with an initial velocity of 30 feet per second. Using the formula H(t) = −16t2 + vt + s, what is the maximum height the soccer ball reaches?

OpenStudy (anonymous):

Maximum height occurs where H '(t) = 0 and H ''(t) < 0. H(t) = -16t^2 + 30t + 0 H '(t) = -32t + 30 0 = -32t + 30 t = 30/32 = 15/16 H ''(t) = -32 < 0 so H(15/16) is a maximum. H(15/16) = -16(15/16)^2 + 30(15/16) + 0 = -225/16 + 450/16 = 225/16 feet

OpenStudy (migitmack):

@doubledrive u copy and paste this from yahoo answers didn't u -.-

OpenStudy (anonymous):

did it help u??

OpenStudy (migitmack):

yeah

OpenStudy (migitmack):

thx

OpenStudy (anonymous):

lol, ur welcoke